OpenStudy (mnr134):

What is the graph of 7x+y=-7

OpenStudy (bri-bri2017):

a good graphing site is desmo if you ever need a graph for an equation. Not much help on how to do it since I don't know myself...

OpenStudy (aveline):

7x+y=-7 It's going to be easiest if you move y to one side of the equation. Do you know how to do this?

OpenStudy (aveline):

Hint: move 7x to the other side

OpenStudy (mnr134):

sorry i put the wrong one

OpenStudy (mnr134):

on the graghs the answers are A.(-7,-1)B.(-1,-7)C.(1,7)D.(7,1)

OpenStudy (aveline):

Your equation is: y=-7x-7 Plug in different x-values to see which one works.

OpenStudy (mnr134):

Whitemonsterbunny17

OpenStudy (mnr134):

no my equation is 7x+y=-7

Directrix (directrix):

If you enter the equation 7x+y=-7 over at Desmos.com you should get a graph that looks like this.

Directrix (directrix):

@mnr134

Directrix (directrix):

This is what you posted as the question: What is the graph of 7x+y=-7

OpenStudy (aveline):

"no my equation is 7x+y=-7" y=-7x-7 is the same thing. Look at the graph directrix posted.

Directrix (directrix):

That is what is graphed. Look at the left of the attachment and you will see that equation.

Directrix (directrix):

I don't know what is the source of confusion about this. Anybody know?

OpenStudy (mnr134):

so whats the answer

OpenStudy (aveline):

._.

OpenStudy (aveline):

Look at the graph Directrix posted

Directrix (directrix):

The graph is the answer to your question.

Directrix (directrix):

This.

Directrix (directrix):

The whole thing. ^^^^

OpenStudy (mnr134):

thank you

Directrix (directrix):

You are welcome.